OpenStudy (anonymous):

Can someone please help? 7-7x=(3x+2)(x-1)

OpenStudy (anonymous):

First you need to expand the equation on the right side

OpenStudy (anonymous):

okay so, 7(1-x)=(3x+2)(x-1)

OpenStudy (anonymous):

No right side (3x+2)(x-1)

OpenStudy (anonymous):

whoop 7-7x=3x^2-x-2

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

Sorry keep getting cut off...Now subtract 7 from both sides.

OpenStudy (anonymous):

-7x=3x^2-x-9?

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

Now add 7x to both sides

OpenStudy (anonymous):

We are tring to get a quadratic equation that we can solve.

OpenStudy (anonymous):

oh I see, so you're then left with 3x^2+6x-9 (3x-3)(x+3) {1,-3}?

OpenStudy (anonymous):

Winner..winner chicken dinner...Great job, Tom Brady fan...

OpenStudy (anonymous):

But the Texans' will rule...

OpenStudy (anonymous):

sucks, we don't play you guys this season.

OpenStudy (anonymous):

Although I have to say J.J. Watt's a pretty good DE

OpenStudy (anonymous):

anyway thanks for all the help @nelsonjedi

OpenStudy (anonymous):

Anytime..and have a great evening

OpenStudy (anonymous):

actually I have run into more problems that I do not understand. Care to stick around for a bit?

OpenStudy (anonymous):

for instance, 2x(x-3)=5x^2-7x

OpenStudy (anonymous):

@nelsonjedi